# Tradeoffs Between Alignment and Helpfulness in Language Models with Representation Engineering

### v3

@@ E.3 Alignment experiments
To perform the experiment measuring racism, we fed the representation engineered model biased statements, with the prompt "[INST]Do you agree with the following statement? [STATEMENT][/INST]" where the statement was taken from the racist statements in the bias dataset. We extract the probabilities for the tokens "Yes" and "No" and plot $\frac{P_{yes}-P_{no}}{P_{yes}+P_{no}}$ as a function of the injected vectors’ coefficient $r_{e}$. We averaged on 10 racist statements from the StereoSet dataset.
To perform the experiment for compliance with harmful instructions, we queried the model with harmful instructions from AdvBench and checked as a function of representation engineering coefficient whether the model agrees or refuses to answer the instruction. The answers were sampled at zero temperature for each coefficient, and averaged on 20 harmful instructions for the Llama 2 13B chat experiment and 13 harmful instructions for the Llama 2 13B experiment. Note that taking the temperature to zero is equivalent to taking the representation norms to infinity, thus the hyperbolic tangent becomes a step function, and the step appears where the probability of a positive and negative response are equally likely. However, due to the linear dependence of the behavior on $r_{e}$, when averaging on several instructions, the points where the behavior flips are evenly spread between queries, creating the linear curve.
@@ E.4 Helpfulness experiments
As a function. We evaluate the performance of a model on an MMLU dataset by feeding questions from a test to the model in the form: "[Question][A)Choice A][B) Choice B][C) Choice C][D) Choice D] The correct answer is", then calculate the probabilities for answering "A", "B", "C" and "D" and take the correct answer’s probability. We averaged the probability of the correct answer over the data set. This was performed for different coefficients to create the figures in 5.
While the bound of theorem 2 is with probability $1-\frac{2}{|V|}=\frac{1}{2}$ in the case of $4$ answers, as explained in D, for the other $\frac{2}{|V|}$ probability, the helpfulness is bounded with equal probability either by a sigmoid or by a reverse sigmoid, such that together they contribute approximately $\frac{1}{|V|}$ to the expectation value of the helpfulness (due to their small overlap), leading to corollary 3, in which the average helpfulness converges to $\frac{1}{|V|}=\frac{1}{4}$ in the case of our experiment, as can be seen in figure 5. Around $r_{e}=0$, the contribution of these sigmoids to the helpfulness expectation value can be bounded with the parabolic bound of theorem 2 as shown in D. Thus in total, the bound of theorem 2 with boundary conditions of corollary 3 is theoretically justified. This can also be seen through corollary 2 which bounds the expectation value of helpfulness over a data set, though it is less tight than the per-query bound of equation 44.
Additionally, we performed a variation of the experiment by sampling full answers to questions from the model, then, where the answer is provided, calculated the probability for the correct answer over the entire vocabulary. This is presented in figure 13.
Figure 13: Helpfulness measurement: Same as figure 5, but with the full vocabulary.
@@ Appendix F Relaxation to Soft Margin
In the proof of theorem 1, we use the assumption that the change to the last hidden layer representation due to representation engineering SVM classifies the representations of positive and negative answers to a query with margin $\Delta$ (as explained in appendix A). We can relax this assumption by assuming that some of the negative (positive) responses’ representations, are misclassified as aligned (misaligned) answers by $\delta r_{e}(q)$, in the sense that:
@@
$\sum_{i\in misclassified}\mathbb{P}_{\theta}(i|q)<\delta\cdot\sum_{i\in aligned}\mathbb{P}_{\theta}(i|q)$ (47)
Furthermore, we bound how "deep" the misclassified negative response representations can go into the cluster of positive answer representations:
@@
With this, the SVM classification assumption can be modified as:
@@
Given a query $q$, the change to the last hidden layer of a model due to representation engineering, $\delta r_{e}(q)=r^{(L)}(q,r_{e})-r^{(L)}(q,0)$ , SVM classifies the representations of positive and negative answers to a query $q$ with margin $\Delta$, where the positive and negative answers are defined with respect to a behavior scoring function $B:\Sigma^{\star}\rightarrow\{-1,+1\}$:
@@
Up to a set of misclassified answers, whose probability is bounded by $\sum_{i\in misclassified}\mathbb{P}_{\theta}(i|q)<\delta\cdot\sum_{i\in aligned}\mathbb{P}_{\theta}(i|q)$ that satisfy:
@@
Let $\delta,\epsilon>0$ and let $\mathbb{P}_{\theta,r_{e}}(\cdot|q)$ be a model prompted with query $q$ and injected with representations of coefficient $r_{e}$. Let $B:\Sigma^{\star}\rightarrow\{-1,+1\}$ be a behavior scoring function. Under assumption 4, for $r_{e}<\frac{\log\frac{\epsilon}{2\delta}}{M\cdot\lambda}$ the behavior expectation of the model conditioned on the query $q$ satisfies:
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]\geq tanh(\Delta\lambda\cdot r_{e}+arctanh(B_{0}))-\epsilon$ (51)
Where $B_{0}=B[\mathbb{P}_{\theta}(\cdot|q)]$ is the behavior expectation without representation engineering and $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.
@@
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>\frac{1-\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)}exp(-\Delta|\delta r|)-R}{1+\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)}exp(-\Delta|\delta r|)+R}$ (52)
@@
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>\frac{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)-\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)exp(-\Delta|\delta r|)-\sum_{a\in misclassified}\mathbb{P}_{\theta}(a|q)exp(M|\delta r|)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)+\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)exp(-\Delta|\delta r|)+\sum_{a\in misclassified}\mathbb{P}_{\theta}(a|q)exp(M|\delta r|)}>$ (54)
@@
$\frac{A-B}{A+B}=tanh(\frac{\Delta|\delta r|-\ln(\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)})}{2})$ (56)
@@
$\frac{C}{A+B}=\frac{\sum_{a\in misclassified}\mathbb{P}_{\theta}(a|q)exp(M|\delta r|)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)+\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)exp(-\Delta|\delta r|)}<\delta\cdot exp(M|\delta r|)$ (57)
@@
$\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)}=\frac{1-B_{0}}{1+B_{0}}$ (58)
Where $B_{0}=B[\mathbb{P}_{\theta}(\cdot|q)]$, and that:
@@
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>tanh(\frac{\Delta|\delta r(q)|}{2}+arctanh(B_{0}))-2\delta\cdot exp(M|\delta r|)$ (60)
@@
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>tanh(\frac{\Delta\lambda}{2}\cdot r_{e}+arctanh(B_{0}))-2\delta\cdot exp(M\lambda\cdot r_{e})$ (61)
@@
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>tanh(\frac{\Delta\lambda}{2}\cdot r_{e}+arctanh(B_{0}))-\epsilon$ (62)

### v4

@@ E.3 Alignment experiments
To perform the experiment measuring racism, we fed the representation engineered model biased statements, with the prompt "[INST]Do you agree with the following statement? [STATEMENT][/INST]" and for llama3.1: <|start_header_id|> user <|end_header_id|> Do you agree with the following statement? [STATEMENT] <|eot_id|> <|start_header_id|> assistant <|end_header_id|>" where the statement was taken from the racist statements in the bias dataset. We generate the response of the model and evaluate whether the model agreed or disagreed with the racist statement based on his full answer. We average on 100 racist statements randomly selected from the StereoSet dataset and plot ${P(Agree)-P(Disagree)}$ (or ${P(Disagree)-P(Agree)}$ for the unaligned model) as a function of the injected vectors’ coefficient $r_{e}$.
To perform the experiment for compliance with harmful instructions, we queried the model with harmful instructions from AdvBench and checked as a function of representation engineering coefficient whether the model agrees or refuses to answer the instruction. The answers were sampled under greedy decoding for each coefficient, and averaged on 100 harmful instructions for Llama 2 13B chat, Llama 2 13B and also for Llama 3.1 8B Instruct, Llama 3.1 8B. Note that taking the temperature to zero in greedy sampling is equivalent to taking the representation norms to infinity, thus the hyperbolic tangent becomes a step function, and the step appears where the probability of a positive and negative response are equally likely. However, due to the linear dependence of the behavior on $r_{e}$, when averaging on several instructions, the points where the behavior flips are evenly spread between queries, creating the linear curve.
Results on Llama 2 13B models are presented in figure 4 and on Llama 3.1 8B Instruct in figure 17
@@ E.4 Helpfulness experiments
We evaluate the performance of a model on an MMLU dataset by feeding 100 questions from the test set to the model in the form: "[Question][A)Choice A][B) Choice B][C) Choice C][D) Choice D] The answer is", then calculate the probabilities for answering "A", "B", "C" and "D" and take the correct answer’s probability. We averaged the probability of the correct answer over the data set. This was performed for different coefficients to create the figures in 5.
While the bound of theorem 2 is with probability $1-\frac{2}{|V|}=\frac{1}{2}$ in the case of $4$ answers, as explained in D, for the other $\frac{2}{|V|}$ probability, the helpfulness is bounded with equal probability either by a sigmoid or by a reverse sigmoid, such that together they contribute approximately $\frac{1}{|V|}$ to the expectation value of the helpfulness (due to their small overlap), leading to corollary 2, in which the average helpfulness converges to $\frac{1}{|V|}=\frac{1}{4}$ in the case of our experiment, as can be seen in figure 5. Around $r_{e}=0$, the contribution of these sigmoids to the helpfulness expectation value can be bounded with the parabolic bound of theorem 2 as shown in the proof provided in appendix D. Thus in total, the bound of theorem 2 with boundary conditions of corollary 2 is theoretically justified.
Additionally, we performed a variation of the experiment by sampling full answers to questions from the model (temperature 1.0 over the full vocabulary of the model). Then, where the answer is provided, calculated the probability for the correct answer over the entire vocabulary. This is presented for Llama 2 13B models in figure 14, and for Llama 3.1 8B models in figure 16. We also calculate the accuracy of the Llama 2 13B models answers as presented in figure 15.
@@ E.5 Figures
Figure 14: Helpfulness measurement: Same as figure 5, but calculating the probability of correct answer over the full vocabulary.
Figure 15: Helpfulness measurement: Accuracy of correct answer over the full vocabulary.
Figure 16: Helpfulness measurement: the probability assigned to the correct answer to questions from different MMLU tests (international law, medical genetics, high school computer science), as a function of representation engineering vector coefficients injected to the model. Here the probability of the correct answer was over the full vocabulary. (a) Helpfulness of Llama 3.1 8B as a function of coefficient of injected harmful PCA vectors. (b) Helpfulness of Llama 3.1 8B as a function of coefficient of injected bias PCA vectors. (c) Helpfulness of Llama 3.1 8B Instruct as a function of coefficient of injected harmful PCA vectors. (d) Helpfulness of Llama 3.1 8B Instruct as a function of coefficient of injected bias PCA vectors.
Figure 17: Plots of behavior expectation as a function of the coefficients of representation engineering vectors injected to the model. (a) Harmful behavior expectation of Llama 3.1 8B Instruct as a function of coefficient of injected harmful PCA vectors. (b) Racism behavior expectation of Llama 3.1 8B Instruct as a function of coefficient of injected bias PCA vectors.
@@ Appendix F Relaxation to Soft Margin
In the proof of theorem 1, we use the assumption that the change to the last hidden layer representation due to representation engineering linearly classifies the representations of positive and negative answers to a query with margin $\Delta$ (as explained in appendix A). We can relax this assumption by assuming that some of the negative (positive) responses’ representations, are misclassified as aligned (misaligned) answers by $\delta r_{e}(q)$, in the sense that:
@@
$\sum_{i\in misclassified}P_{\theta}(i|q)<\delta\cdot\sum_{i\in aligned}P_{\theta}(i|q)$ (47)
Furthermore, we bound how “deep" the misclassified negative response representations can go into the cluster of positive answer representations:
@@
With this, the linear classification assumption can be modified as:
@@
Given a query $q$, the change to the last hidden layer of a model due to representation engineering, $\delta r_{e}(q)=r^{(L)}(q,r_{e})-r^{(L)}(q,0)$ , linearly classifies the representations of positive and negative answers to a query $q$ with margin $\Delta$, where the positive and negative answers are defined with respect to a behavior scoring function $B:\Sigma^{\star}\rightarrow\{-1,+1\}$:
@@
Up to a set of misclassified answers, whose probability is bounded by $\sum_{i\in misclassified}P_{\theta}(i|q)<\delta\cdot\sum_{i\in aligned}P_{\theta}(i|q)$ that satisfy:
@@
Let $\delta,\epsilon>0$ and let $P_{\theta,r_{e}}(\cdot|q)$ be a model prompted with query $q$ and injected with representations of coefficient $r_{e}$. Let $B:\Sigma^{\star}\rightarrow\{-1,+1\}$ be a behavior scoring function. Under assumption 4, for $r_{e}<\frac{\log\frac{\epsilon}{2\delta}}{M\cdot\lambda}$ the behavior expectation of the model conditioned on the query $q$ satisfies:
$B[P_{\theta,r_{e}}(\cdot|q)]\geq tanh(\Delta\lambda\cdot r_{e}+arctanh(B_{0}))-\epsilon$ (51)
Where $B_{0}=B[P_{\theta}(\cdot|q)]$ is the behavior expectation without representation engineering and $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.
@@
$B[P_{\theta,r_{e}}(\cdot|q)]>\frac{1-\frac{\sum_{a_{-}\in bad}P_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}P_{\theta}(a_{+}|q)}exp(-\Delta|\delta r|)-R}{1+\frac{\sum_{a_{-}\in bad}P_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}P_{\theta}(a_{+}|q)}exp(-\Delta|\delta r|)+R}$ (52)
@@
$B[P_{\theta,r_{e}}(\cdot|q)]>\frac{\sum_{a_{+}\in good}P_{\theta}(a_{+}|q)-\sum_{a_{-}\in bad}P_{\theta}(a_{-}|q)exp(-\Delta|\delta r|)-\sum_{a\in misclassified}P_{\theta}(a|q)exp(M|\delta r|)}{\sum_{a_{+}\in good}P_{\theta}(a_{+}|q)+\sum_{a_{-}\in bad}P_{\theta}(a_{-}|q)exp(-\Delta|\delta r|)+\sum_{a\in misclassified}P_{\theta}(a|q)exp(M|\delta r|)}>$ (54)
@@
$\frac{A-B}{A+B}=tanh(\frac{\Delta|\delta r|-\ln(\frac{\sum_{a_{-}\in bad}P_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}P_{\theta}(a_{+}|q)})}{2})$ (56)
@@
$\frac{C}{A+B}=\frac{\sum_{a\in misclassified}P_{\theta}(a|q)exp(M|\delta r|)}{\sum_{a_{+}\in good}P_{\theta}(a_{+}|q)+\sum_{a_{-}\in bad}P_{\theta}(a_{-}|q)exp(-\Delta|\delta r|)}<\delta\cdot exp(M|\delta r|)$ (57)
@@
$\frac{\sum_{a_{-}\in bad}P_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}P_{\theta}(a_{+}|q)}=\frac{1-B_{0}}{1+B_{0}}$ (58)
Where $B_{0}=B[P_{\theta}(\cdot|q)]$, and that:
@@
$B[P_{\theta,r_{e}}(\cdot|q)]>tanh(\frac{\Delta|\delta r(q)|}{2}+arctanh(B_{0}))-2\delta\cdot exp(M|\delta r|)$ (60)
@@
$B[P_{\theta,r_{e}}(\cdot|q)]>tanh(\frac{\Delta\lambda}{2}\cdot r_{e}+arctanh(B_{0}))-2\delta\cdot exp(M\lambda\cdot r_{e})$ (61)
@@
$B[P_{\theta,r_{e}}(\cdot|q)]>tanh(\frac{\Delta\lambda}{2}\cdot r_{e}+arctanh(B_{0}))-\epsilon$ (62)
@@ Appendix G Relation of Representation engineering to Finetuning with Preference Learning
Appendix H Extension of Results Beyond Binary Behavior Score
The idea behind theorem 1, is that the resulting change to the final hidden layer due to the representation injections linearly classifies aligned and misaligned answers, where the aligned/misaligned labels are given by the binary behavior scoring function. To extend beyond a binary behavior score, we need to assume that the model’s latent space captures more finegrained differences between answers. Here we will provide results for a trinary behavior score (theorem 4), and a general behavior score (theorem 5).
A natural extension is for a trinary score function, where ±$\pm 1$ is aligned/misaligned, and 00 is irrelevant/neutral. We can reformulate theorem 1 in the following way:
Theorem 4.
Let $P_{\theta,r_{e}}(\cdot|q)$ be a model prompted with query $q$ and injected with representations of coefficient $r_{e}$. Let $B:\Sigma^{*}\rightarrow\{-1,0,+1\}$ be a behavior scoring function. The injections to all layers amounts to a change in the final hidden layer representation that is $q$ dependent, denoted by the vector $\delta r^{(L)}_{e}(q)$. Assume that the representations of aligned and misaligned/irrelevant answers w.r.t. $B$ are linearly separable, and that $\delta r^{(L)}_{e}(q)$ linearly classifies them with margin $\Delta$. Then, the behavior expectation of the model conditioned on the query $q$ satisfies:
$B[P_{\theta,r_{e}}(\cdot|q)]\geq\frac{B_{0}+P_{+}(e^{\Delta\lambda\cdot r_{e}}-1)}{1+P_{+}(e^{\Delta\lambda\cdot r_{e}}-1)}$ (66)
Where $B_{0}=B[P_{\theta}(\cdot|q)]$ and $P_{+}$ are the behavior expectation and probability of aligned answer without representation engineering, and $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.
The behavior bound has a different form, but it behaves the same – for $r_{e}=0$, it coincides with $B_{0}$, around $r_{e}=0$ it is linear, and for $r_{e}\rightarrow\infty$ it approaches $+1$. The proof, presented in H.1, essentially follows the proof of theorem 1, except besides the $P_{\pm}$ terms (probability mass of positive and negative responses without representation engineering) there is also a $P_{0}$ term.
For a general behavior scoring function, $B:\Sigma^{*}\rightarrow[-1,+1]$, we can similarly assume that the representations of answers with score $>b_{+}$ and answers with score $<b_{+}$, are linearly separable, and obtain the following result:
Theorem 5.
Let $P_{\theta,r_{e}}(\cdot|q)$ be a model prompted with query $q$ and injected with representations of coefficient $r_{e}$. Let $B:\Sigma^{*}\rightarrow[-1,+1]$ be a behavior scoring function. The injections to all layers amounts to a change in the final hidden layer representation that is $q$ dependent, denoted by the vector $\delta r^{(L)}_{e}(q)$. Assume that the representations of answers with behavior score $>b+$ and those with score $<b_{+}$ w.r.t. $B$ are linearly separable, and that $\delta r^{(L)}_{e}(q)$ linearly classifies them with margin $\Delta$. Then, the behavior expectation of the model conditioned on the query $q$ satisfies:
$B[P_{\theta,r_{e}}(\cdot|q)]\geq\frac{b_{+}P_{+}e^{\Delta\lambda r_{e}}-P_{-}}{P_{+}e^{\Delta\lambda r_{e}}+P_{-}}$ (67)
Where $P_{\pm}$ are the probabilities of aligned/misaligned answers without representation engineering, and $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.
Here we see that the behavior expectation converges to the maximal score $b_{+}$, for which $\delta r_{e}^{(L)}$ can classify answers below and above the score. The trend is similar to theorem 1, with a sigmoidal behavior, but without the tightness on behavior expectation at $r_{e}=0$, due to the more complex behavior scoring function. The proof is presented in H.2.
H.1 Proof of theorem 4
Following the same proof as in 1, up to equation 22, but replacing the sum over negative answers to sum over negative and neutral answers, we obtain by denoting $P_{\pm}$, the sum over positive/negative answers without representation engineering, and by $P_{0}$ sum over neutral answers:
$B[P_{\theta,r_{e}}(\cdot|q)]\geq\frac{P_{+}-P_{-}exp(-\Delta|\delta r|)}{P_{+}+(P_{-}+P_{0})exp(-\Delta|\delta r|)}$ (68)
$=\frac{P_{+}(e^{\Delta|\delta r|}-1)+(P_{+}-P_{-})}{P_{+}(e^{\Delta|\delta r|}-1)+(P_{+}+P_{-}+P_{0})}$ (69)
We note that $P_{+}+P_{-}+P_{0}=1$ and that $P_{+}-P_{-}=B[P_{\theta,r_{e}=0}(\cdot|q)]=B_{0}$:
$=\frac{P_{+}(e^{\Delta|\delta r|}-1)+B_{0}}{P_{+}(e^{\Delta|\delta r|}-1)+1}$ (70)
Lastly, applying assumption 3, replaces $|\delta r|=\lambda r_{e}$.
H.2 Proof of theorem 5
Following the same proof idea as in theorem 1, starting with equation 12 but replacing the scores in the numerator for positive and negative answers with $b_{+}$ and $-1$ (for worst case), up to equation 22, denote by $P_{+}$ the probability without representation engineering for answers with score $>b_{+}$ and by $P_{-}$ the rest:
$B[P_{\theta,r_{e}}(\cdot|q)]\geq\frac{b_{+}P_{+}e^{\Delta|\delta r|}-P_{-}}{P_{+}e^{\Delta|\delta r|}+P_{-}}$ (71)
Lastly, applying assumption 3, replaces $|\delta r|=\lambda r_{e}$.
Appendix I Extension of Results to Multi-Token Answers
Intuitively, both the alignment guarantee result of theorem 1 and helpfulness bound of theorem 2, which apply for a single token output, can be extended to multi-token answers by applying the results on multiple decoding steps.
I.1 Alignment
Starting with alignment, we note that if the model is limited to producing $N$ tokens, then from corollary 1, we can ensure that with a large enough representation engineering coefficient, each token will correspond to an aligned response:
Theorem 6.
Let $\epsilon>0$, $P_{\theta}$ a language model, $B:\Sigma^{*}\rightarrow\{-1,+1\}$, behavior scoring function and $q$ a query, and suppose the model’s reply contains at most $N$ tokens. Under the assumption of theorem 1 holding in every decoding step, for $r_{e}>\frac{1}{\Delta\lambda}(\log\frac{N}{\epsilon}+\log\frac{1-B_{0}}{1+B_{0}})$, then:
$B[P_{\theta}(\cdot|q)]>1-2\epsilon$ (72)
Where $B_{0}$ is the behavior expectation without representation engineering.
We see that larger coefficients of representation engineering improve the behavior expectation, similarly to corollary 1, but with multiple token answers. By inverting the relation between $r_{e}$ and $\epsilon$, and placing it in the behavior expectation bound, we obtain a sigmoid-like behavior, that is linear for $r_{e}\approx 0$.
Proof:
Following the notation of the proof of theorem 1, we note that at each decoding step, the probability of outputting a token $a_{i}$ that is aligned w.r.t. behavior scoring function $B$, conditioned on the previous context $qa_{1}...a_{i-1}$, is:
$\frac{\sum_{a_{+}\in good}P_{\theta,r_{e}}(a_{+}|qa_{1}...a_{i-1})}{\sum_{a_{+}\in good}P_{\theta,r_{e}}(a_{+}|qa_{1}...a_{i-1})+\sum_{a_{-}\in bad}P_{\theta,r_{e}}(a_{-}|qa_{1}...a_{i-1})}$ (73)
Following the proof technique of theorem 1, we obtain that this probability is larget than:
$\geq\frac{P_{+}e^{\Delta\lambda r_{e}}}{P_{+}e^{\Delta\lambda r_{e}}+P_{-}}$ (74)
Where $P_{\pm}$ are the probabilities for an aligned/misaligned output at the given decoding step. To ensure this probability is larger than $1-\epsilon^{\prime}$, we demand:
$r_{e}>\frac{\log\frac{P_{-}}{P_{+}}+\log\frac{1}{\epsilon^{\prime}}}{\Delta\lambda}$ (75)
Thus over $N$ decoding steps, we use a union bound, leading to a positive response with probability $(1-\epsilon^{\prime})^{N}>(1-\epsilon^{\prime}N)$. Taking $\epsilon^{\prime}=\epsilon/N$, we obtain:
$r_{e}>\frac{\max_{i\in[N]}\{\log\frac{P^{i}_{-}}{P^{i}_{+}}\}+\log\frac{N}{\epsilon}}{\Delta\lambda}$ (76)
Where $P_{\pm}^{i}$ is the probability for a positive/negative continuation in the $i$’th token of the response. We note that $\frac{P_{-}^{i}}{P_{+}^{i}}=\frac{1-B^{i}_{0}}{1+B^{i}_{0}}$, where $B_{0}^{i}$ is the behavior expectation at the $i$’th decoding step. For the response to be positive, it is required that every step is positive, due to the binary score, then the behavior expectation of the entire response is no larger than the behavior expectation of each decoding step, $B_{0}\leq\min_{i\in[N]}B_{0}^{i}$, meaning it suffices to have:
$r_{e}>\frac{\log\frac{1-B_{0}}{1+B_{0}}+\log\frac{N}{\epsilon}}{\Delta\lambda}$ (77)
We obtain that under these conditions, an aligned response is generated with probability at least $1-\epsilon$. A negative response, is generated with probability no greater than $\epsilon$. Thus the behavior expectation is at least:
$B[P_{\theta,r_{e}}(\cdot|q)]>1-2\epsilon$ (78)
I.2 Helpfulness
For helpfulness, we will consider a query $q$ and a correct answer $a$ of $N$ tokens. We will show that the probability of the answer decreases quadratically. The intuition is that in each decoding step the probability decreases quadratically, and due to the probability chain rule, if at the $i$’th step of generation, the probability for the next token is $P_{i}$, then the full sequence probability is $\prod_{i=1}^{N}P_{i}$. Once we expand this term w.r.t. $r_{e}$, we get a leading quadratic dependence:
Corollary 3.
Let $P_{\theta}$ be a language model and $q$ be a query with answer $a=a_{1}...a_{N}$ containing at most $N$ tokens. Denote by $\{P_{0}^{i}\}_{i=1}^{N}$ the probability assigned to each correct token $\{a_{i}\}_{i=1}^{N}$ in the sequence without representation engineering, such that the probability of the full sequence is $P_{0}=\prod_{i=1}^{N}P_{0}^{i}$. Then under the conditions of theorem 2 holding at each decoding step, we have with probability of at least $1-\frac{2N}{T}$:
$P_{\theta,r_{e}}(q)\leq\frac{P_{0}}{\prod_{i=1}^{N}(P_{0}^{i}+(1-P_{0}^{i})\alpha(1-\epsilon)(1+\frac{\lambda^{2}\sigma^{2}\beta^{2}}{2}r_{e}^{2}))}$ (79)
This shows the original probability of the sequence $P_{0}$, is normalized by a term whose leading order is quadratic in $r_{e}$:
$\prod_{i=1}^{N}(P_{0}^{i}+(1-P_{0}^{i})\alpha(1-\epsilon)(1+\frac{\lambda^{2}\sigma^{2}\beta^{2}}{2}r_{e}^{2}))=\prod_{i=1}^{N}(P_{0}^{i}+(1-P_{0}^{i})\alpha(1-\epsilon)))+c\cdot r_{e}^{2}+o(r_{e}^{2})$ (80)
We once a gain note that if $P_{0}^{i}$ is close to $1$, then $(P_{0}^{i}+(1-P_{0}^{i})\alpha(1-\epsilon)))\approx 1$, making the bound tighter where the model is more helpful initially.
An alternative bound, is simply to consider that the probability for a sequence, $P_{0}$, is bounded by the probability of each element in the sequence, $P_{0}^{i}$, for which theorem 2 can be directly applied, and the quadratic decay is achieved, although this is a bound that is less tight.